OpenStudy (anonymous):

Please help!!! Write the equation of a parabola with a vertex at (-2,1), axis y=1, and latus rectum 4 units long.

OpenStudy (anonymous):

I think this is the first time im hearing "latus rectum"

OpenStudy (anonymous):

yeah what is that? lol

OpenStudy (anonymous):

@Hero @ShailKumar

OpenStudy (anonymous):

okay so i just googled it, \[(x-(-2))^2=4(y-1)\] the latus rectum would be the four part.

OpenStudy (zpupster):

a little added help on parabolas (x + 2)² = 4(y - 1) Compare to this standard form: (x - h)² = 4p(y - k) which has these properties: 1. vertex is the point (h,k) 2. line of symmetry equation is x = h 3. focus is the point (h,k+p) 4. directrix is the line whose equation is y = k-p 5. length of latus rectum = |4p| h = -2, k = 1, 4p = 4, so p = 1 So for this parabola, 1. vertex is the point (h,k) = (-2,1) 2. line of symmetry equation is x = h or x = -2 3. focus is the point (h,k+p) = (-2,1+1) = (-2,2) 4. directrix is the line whose equation is y = k-p or y = 1-1 or y = 0, which is the x-axis. 5. length of latus rectum = |4p| = 4

OpenStudy (anonymous):

ooh ok lol thanks!!!

OpenStudy (anonymous):

The line segment L1L2 perpendicular to the axis of the parabola and passing through the focus is called latus rectum. It can be proved that its length is 4*|p|.